# Insulator phases of Bose–Fermi mixtures induced by intraspecies next-neighbor interactions

### v3

@@ 1 Introduction
Motivated by the latter, we propose an exploration of the extended Bose–Fermi Hubbard model with two-color fermions and scalar bosons at the hardcore limit, including next-neighbor intraspecies interactions. Here we wish to know what new phenomena can emerge from the interplay between long-range interactions and disparity of particle statistics. Moreover, the inclusion of spinful fermions allows the appearance of spin-selective phases, which are also of high interest [66, 67, 68, 69, 70], and the restriction in the form of the hardcore limit keeps a tractable Hilbert space for simulation while also allowing the extrapolation of its results to soft-core models, as shown in previous works [67, 68]. There are only a few studies where these effects are considered for Bose–Fermi mixtures, with one example being an investigation with polarized fermions [87], then we believe this work significantly expands our understanding of this topic. Besides the convergence of on-site and long-range interactions leads us to a multi-parameter model that could be essential to describe physics of recent Bose-Fermi mixtures present in semiconductors heterostructures, where stripes and diverse bidimensional patterns were observed  [33, 34, 35, 36, 37, 38, 39].
@@
Figure 1: Phase diagram as a function of the bosonic $\rho_{\rm{B}}$ and fermionic $\rho_{\rm{F}}$ densities for a balanced $(\rho^{\uparrow}_{\rm{F}}=\rho^{\downarrow}_{\rm{F}})$ Bose–Fermi mixture with intraspecies next-neighbor interactions between only fermions (a) and bosons (b), respectively. Here we consider only rational values for both densities, guaranteeing that the number of carriers is always commensurate with the chain. Each non-vertical line corresponds to the density combinations necessary for the emergence of a particular insulator state, while the rest of the diagram are superfluid phases. The insulators correspond to the well-known mixed Mott insulator (MMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ (dark blue line) and the spin-selective Mott insulator (SSMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}/2=1$ (dark green line), as well as the three phases found: the immiscible charge density wave (ICDW) at $\rho_{\rm{B}}=1/2$ (purple line), the mixed charge density wave (MCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}=0$ (brown line) and the spin-selective charge density wave (SSCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}/2=0$ (light green line). The vertical lines signal special densities that are related to Fig. 2, 4, 6 and 11, here each time these vertical lines intersect with another non-vertical one an insulator phase appears in the corresponding phase diagram (for suitable values of the couplings). The graph only includes part of the range $\rho_{\rm{F}}>1$, since this region corresponds to a reflection of the $\rho_{\rm{F}}<1$ diagram thanks to the particle-hole symmetry of the Hamiltonian shown in Appendix A.
@@ 2 Model
To achieve insight into the physics of this model, we perform a ground-state energy search for a given set of particle fillings $E(N_{\rm B},N_{\rm F}^{\uparrow},N_{\rm F}^{\downarrow})$ and different interaction strengths, using a two-site density matrix renormalization group (DMRG) algorithm based on the MPS language [90] with the help of the TeNPy library [91]. We employ a sweep-dependent maximum bond dimension which starts at a given value between 500 and 1000 and increases every 10 sweeps by 100. The two-site DMRG algorithm truncates the Schmidt coefficients up to a maximum error of $10^{-14}$ or to the maximum bond dimension of the corresponding sweep, nevertheless, since the addition of next-neighbor interactions gives rise to a need for extra degrees of entanglement the latter situation is the general case for our simulations. The stop criteria for the algorithm are given by the energy and entropy errors given by
@@
Here $E_{n}$ and $S_{n}$ are the ground state energy and entanglement entropy at the middle of the lattice at sweep $n$. The denominator in $\Delta E_{n}$ is chosen to avoid numerical overflow if $E_{n}\approx 0$, which is not necessary in the case of entropy. In our program, we set maximum energy and entropy errors of $10^{-5}$ and $10^{-3}$, respectively, then when the wave function has errors above these bounds the optimization stops, which showed to be enough to achieve convergence.
@@ 3 Fermionic next-neighbor interactions
The key findings of this work are shown in the phase diagrams from Fig. 1, where the different insulator phases that can appear at a given filling are summarized. Each diagram assumes that all densities attain rational values, which guarantees the commensurability between the chain and each component density. Moreover, since all density relations that define the possible configurations a certain insulator can have are linear functions of $\rho_{\rm B}$, $\rho_{\rm F}^{\uparrow}$, and $\rho_{\rm F}^{\downarrow}$ (see Sec. 1), each insulator is represented by a straight line in the phase diagram. This schematic also includes the results already known for the single component models: For $\rho_{\rm B}=0$ we have the band insulators at $\rho_{\rm F}=0$ and $2$, and the Mott insulator at half-filling $(\rho_{\rm F}=1)$, well known from the original Fermi-Hubbard model[92]. On the other hand, for $\rho_{\rm F}=0$ we find the analogous band insulator at $\rho_{\rm B}=0$ and the bosonic Mott insulator at unit-filling $(\rho_{\rm B}=1)$ [93, 94, 95, 74, 96]. In the case of $V_{\rm BB}\neq 0$ (Fig. 1(b)) the CDW insulator is found for bosonic chains with next-neighbor interactions at $\rho_{\rm B}=1/2$, which has been found in previous works [76, 50, 77].
@@ 3.1 Two thirds fermionic filling $(\rho_{\rm{F}}=2/3)$
The presence of a non-zero bosonic gap shows the insulating characteristic of this mixture’s component, then it is important to discuss the compressibility of the fermions in the system. Given that $\rho_{\rm B}=1/2$, the conditions for the emergence of the ICDW phase, that is $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$, draw a region in the $\rho^{\downarrow}_{\rm F}-\rho^{\uparrow}_{\rm F}$ plane, inside which we expect no fermionic charge gap since small deviations to the fermionic densities do not affect the conditions for the insulator to emerge. On the other hand, we do expect a fermionic gap on the boundaries, where the transition to the superfluid phases occurs, as it is known to happen for half-filling ($\rho_{B}=\rho_{F}=1/2$) where the state corresponds to the MMI [68], as an example. This is reminiscent of the supersolid phase where crystalline and superfluid orders are simultaneously present [72, 73], nevertheless a further characterization falls outside the scope of this work.
@@ 3.2 Two fifths fermionic filling $(\rho_{\rm{F}}=2/5)$
The emergence of the MCDW phase requires nonzero next-neighbor interactions and a balance of the carrier densities (bosons and fermions), which suggest the formation of Bose–Fermi composites to establish this peculiar insulator state. Another interesting fact is that this incompressible phase can only appear for $\rho_{\rm{B}}+\rho_{\rm{F}}\leq 1$, meaning that all fermions must cooperate to form this unique insulator. The limit case when $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ corresponds to the mutual half-filling situation $\rho_{\rm{B}}=\rho_{\rm{F}}=1/2$ where both orders attain a two-site periodicity, turning into the ICDW as $\mathcal{O}^{\rm MCDW}_{\rho=1/2}=\mathcal{O}^{\rm ICDW}$ from (9) and (10). The large unit cell of this CDW insulator and the fact that the corresponding densities are in general rational (contrary to the MMI and SSMI where the sum of two densities is an integer and the ICDW where the unit cell size is two as expected from a usual CDW phase) shows similarity to ground states found in frustrated magnetism models characterized by rational magnetization plateaus [97, 98, 99]. In this sense, the high nearest-neighbor interaction could be interpreted as a source of frustration that will reduce the effective kinetic energy of the carriers, thus enhancing the stability of the non-trivial CDW insulator (see Fig. 5(a)). While this topic may deviate from the primary focus, it offers a promising avenue for future research on the characterization of this insulator.

### v4

@@ 1 Introduction
Motivated by the latter, we propose an exploration of the extended Bose–Fermi Hubbard model with two-color fermions and scalar bosons at the hardcore limit, including next-neighbor intraspecies interactions. Here we wish to discover what new phenomena can emerge from the interplay between long-range interactions and disparity of particle statistics. Moreover, the inclusion of spinful fermions allows the appearance of spin-selective phases, which are also of high interest [66, 67, 68, 69, 70], and the restriction in the form of the hardcore limit keeps a tractable Hilbert space for simulation while also allowing the extrapolation of its results to soft-core models, as shown in previous articles [67, 68]. There are only a few studies where these effects are considered for Bose–Fermi mixtures, with one example being an investigation with polarized fermions [87], so we believe the present study will significantly expand our understanding of this topic. Furthermore, the convergence of on-site and long-range interactions leads us to a multi-parameter model that could be essential for describing the physics of recent Bose-Fermi mixtures present in semiconductor heterostructures, where stripes and diverse bidimensional patterns have been observed [33, 34, 35, 36, 37, 38, 39].
@@
Figure 1: Phase diagram as a function of the bosonic $\rho_{\rm{B}}$ and fermionic $\rho_{\rm{F}}$ densities for a balanced $(\rho^{\uparrow}_{\rm{F}}=\rho^{\downarrow}_{\rm{F}})$ Bose–Fermi mixture with intraspecies next-neighbor interactions between only fermions (a) and bosons (b), respectively. Here we consider only rational values for both densities, guaranteeing that the number of carriers is always commensurate with the chain. Each non-vertical line corresponds to the density combinations necessary for the emergence of a particular insulator state, while the rest of the diagram shows a superfluid phase. The insulators correspond to the well-known mixed Mott insulator (MMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ (dark blue line) and the spin-selective Mott insulator (SSMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}/2=1$ (dark green line), as well as the three phases found: the immiscible charge density wave (ICDW) at $\rho_{\rm{B}}=1/2$ (purple line), the mixed charge density wave (MCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}=0$ (brown line) and the spin-selective charge density wave (SSCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}/2=0$ (light green line). The vertical lines signal special densities that are related to Fig. 2, 4, 6 and 11, here each time these vertical lines intersect with another non-vertical one an insulator phase appears in the corresponding phase diagram (for suitable values of the couplings). The graph only includes part of the range $\rho_{\rm{F}}>1$, since this region corresponds to a reflection of the $\rho_{\rm{F}}<1$ diagram thanks to the particle-hole symmetry of the Hamiltonian shown in Appendix A.
@@ 2 Model
To achieve insight into the physics of this model, we perform a ground-state energy search for a given set of particle fillings $E(N_{\rm B},N_{\rm F}^{\uparrow},N_{\rm F}^{\downarrow})$ and different interaction strengths, using a two-site density matrix renormalization group (DMRG) algorithm based on the MPS language [90] with the help of the TeNPy library [91]. We employ a sweep-dependent maximum bond dimension that starts at a given value between 500 and 1000 and increases every 10 sweeps by 100. The two-site DMRG algorithm truncates the Schmidt coefficients up to a maximum error of $10^{-14}$ or to the maximum bond dimension of the corresponding sweep; nevertheless, since the addition of next-neighbor interactions gives rise to a need for extra degrees of entanglement, the latter situation is the general case for our simulations. The stop criteria for the algorithm are determined by the energy and entropy errors, given by
@@
Here, $E_{n}$ and $S_{n}$ are the ground-state energy and the entanglement entropy at the middle of the lattice at sweep $n$. The denominator in $\Delta E_{n}$ is chosen to avoid numerical overflow if $E_{n}\approx 0$, which is not necessary in the case of entropy. In our program, we set maximum energy and entropy errors of $10^{-5}$ and $10^{-3}$, respectively, so when the wave function has errors above these bounds, the optimization stops, which proved to be enough to achieve convergence.
@@ 3 Fermionic next-neighbor interactions
The key findings of this work are shown in the phase diagrams from Fig. 1, where the different insulator phases that can appear at a given filling are summarized. Each diagram assumes that all densities attain rational values, which guarantees the commensurability between the chain and each component density. Moreover, since all density relations that define the possible configurations that a certain insulator can have are linear functions of $\rho_{\rm B}$, $\rho_{\rm F}^{\uparrow}$, and $\rho_{\rm F}^{\downarrow}$ (see Sec. 1), each insulator is represented by a straight line in the phase diagram. This schematic also includes the results already known for the single component models: For $\rho_{\rm B}=0$ we have the band insulators at $\rho_{\rm F}=0$ and $2$, and the Mott insulator at half-filling $(\rho_{\rm F}=1)$, well known from the original Fermi-Hubbard model[92]. On the other hand, for $\rho_{\rm F}=0$ we find the analogous band insulator at $\rho_{\rm B}=0$ and the bosonic Mott insulator at unit-filling $(\rho_{\rm B}=1)$ [93, 94, 95, 74, 96]. In the case of $V_{\rm BB}\neq 0$ (Fig. 1(b)) the CDW insulator is found for bosonic chains with next-neighbor interactions at $\rho_{\rm B}=1/2$, which has been found in previous works [76, 50, 77].
@@ 3.1 Two thirds fermionic filling $(\rho_{\rm{F}}=2/3)$
The presence of a non-zero bosonic gap shows the insulating characteristic of this mixture’s component, so it is important to discuss the compressibility of the fermions in the system. Given that $\rho_{\rm B}=1/2$, the conditions for the emergence of the ICDW phase, that is $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$, delineate a region in the $\rho^{\downarrow}_{\rm F}-\rho^{\uparrow}_{\rm F}$ plane, inside of which we expect there to be no fermionic charge gap, since small deviations in the fermionic densities do not affect the conditions for the insulator to emerge. On the other hand, we do expect a fermionic gap at the boundaries, where the transition to the superfluid phases occurs, as is known to happen for half-filling ($\rho_{B}=\rho_{F}=1/2$), where the state corresponds to the MMI [68], as an example. This is reminiscent of the supersolid phase, where crystalline and superfluid orders are present simultaneously [72, 73]. Nevertheless, a further characterization falls outside the scope of this study.
@@ 3.2 Two fifths fermionic filling $(\rho_{\rm{F}}=2/5)$
The emergence of the MCDW phase requires nonzero next-neighbor interactions and a balance of the carrier densities (bosons and fermions), which suggest the formation of Bose–Fermi composites to establish this peculiar insulator state. Another interesting fact is that this incompressible phase can only appear for $\rho_{\rm{B}}+\rho_{\rm{F}}\leq 1$, meaning that all fermions must cooperate to form this unique insulator. The limit case when $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ corresponds to the mutual half-filling situation $\rho_{\rm{B}}=\rho_{\rm{F}}=1/2$ where both orders attain a two-site periodicity, turning into the ICDW as $\mathcal{O}^{\rm MCDW}_{\rho=1/2}=\mathcal{O}^{\rm ICDW}$ from (9) and (10). The large unit cell of this CDW insulator and the fact that the corresponding densities are in general rational (in contrast to the MMI and SSMI, where the sum of two densities is an integer, and the ICDW, where the unit cell size is two, as expected from a usual CDW phase) shows similarity to ground states found in frustrated magnetism models characterized by rational magnetization plateaus [97, 98, 99]. In this sense, the high nearest-neighbor interaction could be interpreted as a source of frustration that will reduce the effective kinetic energy of the carriers, thus enhancing the stability of the non-trivial CDW insulator (see Fig. 5(a)). While this topic may deviate from the primary focus, it offers a promising avenue for future research on the characterization of this insulator.